OpenStudy (anonymous):

HELP !!!!!!! The number 7245 what numbers is it divisible by 2,3,4,5,6,8,9,10,12

OpenStudy (anonymous):

I don't get your question. Well are you asking if the number 7245 is divisible by all of those? :))

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Okay, so to determin if a number is divisible by 2, it has to be even. Since 7245 is not even, it isn't divisible by 2.

OpenStudy (anonymous):

For it to be divisible by 3, the numbers in 7245 should amount to a number divisible by 3 when added. 7+2+4+5= 18 = 1+8 = 9 9 is divisible by 3, so the number 7245 is divisible by 3.

OpenStudy (anonymous):

For it to be divisible by 4, the last 2 numbers should be divisible by 4. 7245 = 45 = Since 45 is not divisible by 4, the number 7245 isn't divisible by 4.

OpenStudy (anonymous):

Numbers ending in 5 or 0 are always divisible by 5. So 7245 is divisible by 5.

OpenStudy (anonymous):

For a number to be divisible by 6, it has to be both divisible by 2 and 3. Since 7245 isn't divisible by 2 but divisible by 3, the number is not divisible by 6.

OpenStudy (anonymous):

For it to be divisible by 8, the last three numbers must be divisible by 8 7245 = 245 = Since 245 is not divisible by 8. 7245 is not divisible by 8

OpenStudy (anonymous):

Just like 3, a number is divisible by 9 if the sum of all it's digits is divisible by 9 7245 = 7+2+4+5 = 18 Since 18 is divisible by 9, the whole number is divisible by 9

OpenStudy (anonymous):

Any number ending in 0 is divisible by 10. Since 7245 doesn't end with a 0, it is not divisible by 10.

OpenStudy (anonymous):

If the number is divisible by both 3 and 4, then the number is divisible by 12.

OpenStudy (anonymous):

Hoped it helped! =))

OpenStudy (anonymous):

Thanks Alot What about the number 21,408

OpenStudy (anonymous):

Just test it with everything I typed and you'll be alright :>